OpenStudy (anonymous):

11/4y+2(-y-6)+11=-67 neeed help working out multi step equations with fractions....

OpenStudy (compassionate):

What do you think the first step will be?

OpenStudy (anonymous):

idk thats where i'm stuck , do i do the fraction first or what

OpenStudy (compassionate):

You can to distribute first. 11/4y+2(-y-6)+11=-67 \[\frac{ 11 }{ 4y } + 2y - 12 + 11 = -67\]

OpenStudy (compassionate):

Can you do the next step?

OpenStudy (anonymous):

no I cant get that part , but do i multiply everything by the denominator to get 11 by itself or do i bring 11/-4 to -67 ?

OpenStudy (compassionate):

No, you want to subtract 11 from both sides, then add 12 to both sides.

OpenStudy (anonymous):

why would u do that if you can subtract -12 from 11 since they are already on the same side which gives u 11/4y-2y-1=-67, then u would add 1 to get 11/4y-2y=-66

OpenStudy (compassionate):

No need for that. Just follow the steps I give you and subtract 11 and add 12

OpenStudy (anonymous):

so ur way and how i did it give u -68, so 11\4y-2y=-68, so do i multiply the y's by 4 to get rid of the fraction?

OpenStudy (compassionate):

You want to multiple by 11 to get rid of fractions.

OpenStudy (anonymous):

now i'm lost i thought u multiplied by the denominator

OpenStudy (compassionate):

Let me type it out. Give me a minute.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what answer did u get ?